OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2

OpenStudy (compassionate):

x = 5

OpenStudy (anonymous):

3x - 3 - x -5 =2 2x-8 = 2 2x =10 x =10/2 = 5